OpenStudy (scarlett_feathers):

I don't do well with word problems......at least I know how to work the fan and medal buttons :D The length of a rectangle is 3 m less than twice the width, and the area of the rectangle is 14 m^2 Find the dimensions of the rectangle.

OpenStudy (mjdennis):

Call the width of the rectangle W. How would you write "twice the width"? How would you write "twice the width minus 3 meters"? What does the problem say about "twice the width minus 3 meters"?

OpenStudy (shaik0124):

rect length=2 *width-3 area=14 l*w=14

OpenStudy (mjdennis):

Hey, @shaik0124 , let's see if we can help him get _better_ at word problems?

OpenStudy (shaik0124):

k

OpenStudy (mjdennis):

(or her. Oops)

OpenStudy (shaik0124):

in word problems please first note down all the details given in the question so that u will be able to solve the problem very easily

OpenStudy (shaik0124):

then try to link the equations u get to find a particular value and subsitute the obtained value wherever required to reach ur final answer

OpenStudy (shaik0124):

scarlett did u get ur answer

OpenStudy (scarlett_feathers):

no. I am not sure how the equation is suppose to be set up. is it 3-2w=14

OpenStudy (shaik0124):

it is given that 3 less than 2 width it's like 3<2w so 2w-3

OpenStudy (scarlett_feathers):

so I reversed it?

OpenStudy (scarlett_feathers):

Can you walk me through how to solve this?

OpenStudy (shaik0124):

so l=2w-3 and given area=14 so area of rectangle is 1*w=14 now u have two equations from two equations u will find l and w

OpenStudy (scarlett_feathers):

\[1*w =14\] \[\frac{ 14 }{ 1}=w ?\]

OpenStudy (shaik0124):

yes now substitute w in the first equation

OpenStudy (scarlett_feathers):

\[l=2\left( 14 \right)-3\] \[l= 28 -3\] \[l= 25\] ?

OpenStudy (shaik0124):

or u can do like this 1*w=14 2*7=14 here l is 2 and w is 7 now subsitute 2 and 7 in first equation if left hand side is equal to right hand side the values of l and w is 2 and 7

OpenStudy (shaik0124):

value of w is 14/l but u substituted on 14 so its wrong

OpenStudy (shaik0124):

its 14/L not 14/1

OpenStudy (shaik0124):